OpenStudy (anonymous):

Is the slope of a line through points (4,6) and (-2,6) zero or undefined?

jimthompson5910 (jim_thompson5910):

notice how the y coordinates are both 6 so you'll have y2 - y1 = 6-6 = 0 in the numerator 0 over any number (but 0 itself) is always 0

jimthompson5910 (jim_thompson5910):

I'm using the slope formula m = (y2 - y1)/(x2 - x1)

OpenStudy (karatechopper):

y1-y2/x1-x2=m m=slope 1/0 will always be undefined 0/1 is always zero.

OpenStudy (anonymous):

Thanks guys! I knew how to find the slope I just didn't know how to tell if it was zero or undefined!